Consider a classical economy with the following characteristic:

Investment function: I = i0 - i1r + i2IC where IC represents investor confidence and r interest rates.

Consumption function: C = a +b(Y-T) - cr

All model parameters are positive.

Suppose that the investor confidence (IC) has increased. How does this change in investor confidence affect the Classical Model? Be sure to mention what happens to (Y, N, P, r, C, S, Private Saving, Public Saving).

Note: Assume that the change in investment does not impact the capital stock. Written discussion and/or graphs are needed.
Answer #1

We know that Aggregate Demand and supply functions at equilibrium can be expressed as

Y = C+ I

So, in the above case we are given consumption and investment functions which we use, to deduce the equilibrium output level.

I= 1-;8+ izle c = atbly-T)-cr Y = o-;8+ / C + a+by-I)-ed y = otizic ta -bT-litc) 1-6 CScanned with CamScanner

We have,

Y =( 1/1-b )*(i_{0} +i_{2} IC + a- bT- (i_{1}+c)r )

When Investment confidence increases, the parameter i2 increases. As a result, Output level at equilibrium increases.

This can also be explained using the IS-LM and AS-AD models.

Increase in investment confidence leads to an upward shift of the IS curve due to increase in Autonomous investmenet expenditure. Due to this shift, more and more investments would occur in many sectors, thereby increasing production capacities. Increased investments also mean increased money demand at a given level of money supply. Hence interest rates also rise. In conclusion, output and interest rate at equilibrium level rises.

- LM il ㅏ B CScanned with CamScanner yl

A shift in the IS also causes the Aggregate Demand to shift upwards. As a result Price level also rises at equilibrium.

e AS 61 Т. Р Ар +y Y Scanned with CamScanner YI

A rise in investment confidence acts as a stimulator for private savings .Since increased private savings means increased investments, there willbe more and more private firms to cater to the public's demand and competition between them would result in acheiving quality at reasonable price levels. So government starts to withdraw its presence in many sectors where private agencies flourish. As a result, public savings decline.